# From Actions to Words: Towards Abstractive-Textual Policy Summarization in RL

### v2

@@
Abstract
Policies generated by Reinforcement Learning (RL) algorithms are difficult to explain to users, as they emerge from the interaction of complex reward structures and neural network representations. Consequently, analyzing and predicting agent behavior can be challenging, undermining user trust in real-world applications. To facilitate user understanding, current methods for global policy summarization typically rely on videos that demonstrate agent behavior in a subset of world states. However, users can only watch a limited number of demonstrations, constraining their understanding. Moreover, these methods place the burden of interpretation on users by presenting raw behaviors rather than synthesizing them into coherent patterns. To resolve these issues, we introduce SySLLM (Synthesized Summary using Large Language Models), advocating for a new paradigm of abstractive-textual policy explanations. By leveraging Large Language Models (LLMs)—which possess extensive world knowledge and pattern synthesis capabilities—SySLLM generates textual summaries that provide structured and comprehensible explanations of agent policies. SySLLM demonstrates that LLMs can interpret spatio-temporally structured descriptions of state-action trajectories from an RL agent and generate valuable policy insights in a zero-shot setting, without any prior knowledge or fine-tuning. Our evaluation shows that SySLLM captures key insights, such as goal preferences and exploration strategies, that were also identified by human experts. Furthermore, in a large-scale user study (with 200 participants), SySLLM summaries were preferred over demonstration-based summaries (HIGHLIGHTS) by a clear majority (75.5%) of participants.
journal: Artificial Intelligence
@@ Abstract
1 Introduction
Reinforcement learning (RL) agents are increasingly deployed in high-stakes domains, yet their opaque decision-making processes remain a barrier to trust and adoption. Explainable RL (XRL) methods aim to bridge this gap by surfacing interpretable accounts of agent behavior. Despite this promise, current XRL methods face significant limitations. Saliency maps [1, 2, 3, 4] highlight critical states but offload pattern discovery to users. Trajectory demonstrations [5, 6, 7, 8] constrain the understanding to cherry-picked scenarios. Rule extraction methods [9, 10, 11, 12], while providing structured representations, often struggle to capture the full complexity of learned policies. Many of these approaches share a common challenge: they present transformed observations rather than synthesized insights, requiring users to manually extrapolate agent intent, adaptability, and failure modes from fragmented evidence [13]. Recent advances in large language models (LLMs) offer a transformative alternative. With their capacity for abstraction, commonsense reasoning, and natural language generation [14], LLMs hold unique potential to synthesize agent trajectories into human-interpretable narratives. However, this promise comes with a challenge, LLMs’ static text-based training contrasts with the dynamic, spatiotemporal nature of reinforcement learning environments [15]. Early applications of LLMs in RL, such as world modeling [16], hint at their promise, but generating faithful global policy summaries—cohesive accounts that generalize beyond isolated states—remains unexplored.
Figure 1: Collecting the experience dataset (Section 4.3).
In this work, we advocate for an abstractive-textual approach for policy summarization. We introduce SySLLM, a framework that leverages LLMs to generate holistic summaries of RL policies by systematically distilling agent trajectories into structured narratives. SySLLM addresses the limitations of previous methods through a two-stage process: (1) converting raw agent-environment interactions into temporally grounded language descriptions (Fig. 1), and (2) prompting LLMs to synthesize these into coherent summaries that capture behavioral patterns, strategic tendencies, and environment-specific adaptations (Fig. 2). SySLLM performs impressively well—it captures agent preferences (as induced by different reward functions), identifies sub-optimal behavior in under-trained agents, and distinguishes between diverse exploration strategies. Beyond summarizing the agent’s behavior, SySLLM offers insights into the effectiveness of the agent’s strategy.
SySLLM demonstrates that LLMs can interpret spatiotemporally structured descriptions of state-action trajectories from an RL agent. While these descriptions are written in natural language, they differ significantly from the kinds of text LLMs are typically trained on, such as stories or dialogue. Instead, they represent a distinct modality: sequential, decision-driven behavior grounded in the dynamics of an environment. Despite this, SySLLM shows that LLMs can process such unfamiliar input and, in a zero-shot setting—without any prior knowledge or fine-tuning—generate valuable policy insights.
Our framework is validated through multi-faceted evaluation. First, comparisons with RL experts reveal that SySLLM summaries align closely with expert annotations, capturing nuanced behaviors that demonstration-based methods do not explicitly highlight. Second, a large-scale user study shows that participants prefer SySLLM over state-of-the-art visual summaries such as HIGHLIGHTS [17], while also achieving equal or better accuracy in agent identification tasks.
In summary, our main contributions are the following:
  • We introduce SySLLM, a framework that takes a step toward generating abstractive-textual explanations of RL policies using LLMs. SySLLM synthesizes structured narratives that reveal agent intent, adaptability, and limitations by systematically analyzing agent trajectories, enabling users to grasp these behaviors without manual pattern matching. SySLLM eases the burden of interpreting agent behavior by digesting complex dynamics into coherent summaries. To the best of our knowledge, this is the first work to produce global, abstractive explanations of an RL agent’s policy in natural language.
  • Generated summaries align with RL experts. We evaluate the effectiveness of SySLLM by comparing its summaries with those written by RL experts, demonstrating substantial alignment and accurate coverage of key policy details. To evaluate faithfulness—an area often prone to issues in automatic summarization [18, 19]—we additionally compute precision scores. Our results show that SySLLM maintains high precision overall, indicating that its summaries rarely introduce unsupported claims.
  • Users prefer our textual summaries. We conduct a large-scale user study with 200 participants showing that participants clearly prefer SySLLM over a leading demonstration-based policy summarization method on a summary preference task (75.5%), while performing equally well or better in an agent identification task.
Figure 2: Generating natural language global policy summaries (Section 4.4).
2 Background and Related Work
2.1 Explainable Reinforcement Learning
Explainable reinforcement learning (XRL) aims to make the behavior of RL agents more understandable to human users. Traditional XRL techniques fall into several major categories: saliency-based methods, demonstration-based summaries, surrogate models, and causal explanations [20].
Saliency and Visualization Techniques
Early XRL work often adopted visualization tools such as saliency maps [1, 2, 3, 4] to indicate which parts of the input space influenced the agent’s decisions. While these tools provide some insight, they require users to manually interpret the spatial and temporal relevance of highlighted regions, leading to potential misinterpretation. Studies show these methods are local and fragile [21], providing little insight into the long-term policy structure.
Demonstration-Based Summaries
Another common approach involves selecting representative episodes or critical states that demonstrate agent behavior [5, 6, 7, 8]. These methods allow users to observe behavior directly, but rely heavily on cherry-picked scenarios, with high cognitive load placed on the user to infer strategic intent from limited data. Work by [22] shows that users often struggle to generalize from such demonstrations, particularly when the agent behavior is stochastic.
Surrogate Models
An alternative strategy uses model distillation [23] to learn interpretable approximations of agent behavior, such as decision trees [24, 25]. These methods strike a balance between interpretability and fidelity: overly simple models fail to capture nuances of the agent’s strategy, while complex surrogates become as opaque as the original policy. Recent works explore bounding MDPs [26] or abstract policy graphs [27] to improve this trade-off, but scalability remains an issue.
Causal Explanations
A complementary direction models the agent’s reasoning through structural causal models or influence graphs [28, 29]. These allow users to query “why” an agent took a specific action and offer contrastive explanations that align with how humans seek understanding. However, such models often require significant prior knowledge of the domain or agent internals, limiting their applicability.
Reward Decomposition
Other methods explain agent behavior by breaking down the overall reward into meaningful components, such as safety, efficiency, or task-specific objectives. This allows explanations that clarify how each reward type influences the agent’s decisions and highlights trade-offs between competing objectives [30, 31, 32]. These approaches help users understand agent priorities and decision rationale by making the contribution of each reward component explicit.
2.2 Large Language Models in Reinforcement Learning
Pre-trained Large Language Models (LLMs) have revolutionized NLP through their ability to distill patterns from vast datasets into coherent text. Autoregressive models like GPT [33] generate human-like narratives by predicting token sequences, while encoder-decoder architectures excel at summarizing complex inputs [34]. Those recent advances have introduced new possibilities for interacting with and explaining RL agents. While much prior work integrates LLMs into the training process e.g., guiding exploration [35] or building world models [16], less attention has been paid to their potential as explanation generators. [36] explore using LLMs to generate local post-hoc explanations of individual decisions based on distilled behavior representations, but such approaches do not aim to provide global summaries of agent behavior.
Trajectory Annotation and Feedback
Some works use LLMs to narrate agent behavior in real-time or provide textual feedback. [37, 38] demonstrate that LLMs can be prompted to interpret actions or suggest improvements mid-episode. These studies show promise for LLM-based commentary, though they often lack grounding in actual agent dynamics, leading to inconsistencies.
Model-Based Reasoning and Simulation
In world modeling applications, LLMs have been used to generate or refine symbolic models of environment dynamics [39]. Here, the LLM acts as a high-level planner or simulator, encoding causal knowledge that guides behavior. This further validates the hypothesis that LLMs can represent abstract behavioral structures, which is essential for high-level summarization.
These applications largely treat LLMs as agents or environment models, not as explainers. When LLMs are used for explanation, they must be carefully grounded in observed experience to avoid plausible-sounding but inaccurate generalizations [40].
2.3 Natural Language Explanations for RL Agents
The use of natural language as a modality for RL explanations offers distinct advantages over visual or rule-based approaches. Language allows for abstraction, temporal coherence, and commonsense reasoning—all of which are challenging to express in saliency maps or decision trees.
Template-Based Explanations
Early works, such as [41], demonstrated that combining learned MDP structures with templated natural language generation could enable contrastive explanations (e.g., “Why didn’t you do X?”) through predicate-based state abstraction, allowing generalization across policy architectures. Later, simpler approaches reemerged, such as [10], which employed rigid, static templates to directly translate policies into interpretable statements (e.g., “If near wall, turn left”), prioritizing accessibility but inheriting brittleness and limited expressivity. More recent systems like [42] also rely on predefined action-condition mappings, facing similar limitations.
Rationalization and Free-Form Generation
Subsequent work adopted neural methods to generate textual rationales. [43] introduced AI rationalization training sequence-to-sequence models to translate trajectories into natural language. Later, [44] introduced graph-based policy summaries that enhance user understanding by clustering states and abstracting the state space.
While those natural language explanations methods offer interpretability, they are inherently limited by their reliance on handcrafted rules and rigid structures, making them brittle and unable to express nuanced or unexpected behaviors. Rationalization methods, which translate individual trajectories into free-form text, tend to focus on local justifications rather than providing a holistic view of agent behavior. These approaches often lack scalability across diverse scenarios and struggle to generalize patterns observed over many episodes. Moreover, they typically do not leverage broader world knowledge or incorporate high-level reasoning, which can limit the clarity, depth, and utility of the resulting explanations for end users.
3 Motivation for Abstractive-Textual Explanations
Despite recent advances in explainable reinforcement learning (XRL), existing methods face a persistent trade-off between expressiveness, scalability, and accessibility. Extractive approaches—such as saliency maps or curated demonstrations—highlight fragments of agent behavior but often fail to integrate them into coherent, high-level narratives. Symbolic models, including logical rules and decision trees, offer structural clarity but require handcrafted features and often struggle to scale in complex or high-dimensional settings.
These limitations point to a need for explanation methods that can synthesize global behavioral insights—capturing patterns, adaptation, and strategy over time—without relying on structured input representations or human-crafted abstractions.
We advocate for a textual-abstractive paradigm in XRL, where natural language is used to summarize agent behavior across episodes. Text offers a flexible medium for conveying temporal dependencies, recurring behaviors, and decision-making logic in a format accessible to non-experts. Unlike extractive or symbolic methods, textual explanations can operate over raw behavioral data without predefined rules or model internals, making them broadly applicable across domains.
Recent advances in large language models (LLMs) open new opportunities for explainability in reinforcement learning. With SySLLM, we introduce a textual-abstractive paradigm that uses natural language to summarize agent behavior over time—capturing patterns, strategic intent, and adaptation. In this work, we show that LLMs are particularly well-suited for this approach, when guided by an appropriate framework, they can recognize sequential structure and express it in fluent, context-aware language. We view this paradigm as a holistic and scalable complement to existing XRL methods, and as part of a broader shift toward language-based interfaces for interpreting learned behavior.
4 SySLLM Method
Our SySLLM (Synthesized Summary using LLMs) method consists of several key components, which we formalize in Algorithm 1. In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language. These captioned experiences are logged into an experience dataset (Sections 4.2 and 4.3). In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.4). To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
4.1 Problem Statement
We define a partially observed Markov Decision Process (POMDP) as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle$, where observations $o\in\mathcal{O}$ are derived from states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ via $O(o|s,a)$. The dynamics are governed by the transition function $T(s^{\prime}|s,a)$, with $R$ as the reward function and $\gamma$ as the discount factor. POMDPs generalize Markov Decision Processes (MDPs), where observations fully reveal the state. The agent operates under a stochastic policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$, mapping each observation $o\in\mathcal{O}$ to a probability distribution over actions $\mathcal{A}$. The objective of the policy is to maximize the expected cumulative discounted reward. Given a POMDP environment and the agent’s policy, our goal is to generate a natural language summary that encapsulates the agent’s strategy and decision-making process. This summary explains the policy’s operational dynamics, effectiveness, and behavior across scenarios. Formally, a summarization is a mapping from $(\mathcal{M},\Pi)$ to $\mathcal{T}$, where $(\mathcal{M},\Pi)$ represents environment-policy pairs, and $\mathcal{T}$ is the set of natural language summaries.
4.2 Captioner
As will be detailed in Section 4.3, our approach utilizes captioners to leverage LLMs for summarizing an agent’s policy. We convert each observation $o\in\mathcal{O}$ and action $a\in\mathcal{A}$ into natural language descriptions that highlight their key features. In this subsection, we formalize the role and function of the captioners.
Observation Captioner:
The observation captioner, $C_{\text{observation}}$, translates observations into natural language descriptions:
$C_{\text{observation}}:\mathcal{O}\rightarrow\Sigma^{\star}.$
Where $\Sigma^{\star}$ represents the set of all possible strings. This function distills complex state information into comprehensible text, such as translating a grid map into a description of object locations.
Action Captioner:
Similarly, the action captioner, $C_{\text{action}}$, maps actions to natural language representations:
$C_{\text{action}}:\mathcal{A}\rightarrow\Sigma^{\star}.$
Similar to past work [45, 46] we assume the availability of a captioner language annotator. The process of generating captions is supported by a range of established methods, which are further discussed in Section 8.
4.3 Collecting the Agent’s Experience Dataset
We describe the first phase of Algorithm 1, detailed in Lines 114. We assume that the environment ($ENV$) supports two functions. The first function is RESET, which resets the environment to its initial state and returns the initial observation, and STEP, which takes an action as input and returns the next observation, the reward for that action, a boolean indicating whether the episode has ended, and additional information. First, we initialize the experience dataset $ED$ (Line 1). The for loop (Line 2) iterates over multiple episodes, initializing the step counter $t=1$ at the beginning of each episode (Line 3) and obtaining the initial observation $o_{1}$ (Line 4). The variables $epReward$ and donedone are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56). The while loop (Line 7) continues until the episode concludes. During each iteration of the while loop, an action is sampled from the policy $\pi$ (Line 8), and both the observation and action are converted into natural language representations by the captioners. These representations are added to the experience dataset along with the cumulative reward (Line 9). Subsequently, the environment is updated using the selected action, and the next observation is retrieved (Line 10). Finally, the cumulative reward and step index are updated accordingly (Lines 1112). The information stored in the experience dataset is detailed in B.
4.4 Generating Natural Language Summaries
We move on to addressing the second phase of Algorithm 1, covered in Lines 1517. To generate natural language summaries, we use the experience dataset and an LLM. The process begins by formatting the experience dataset into a structured format suitable for integration into a prompt. This format captures the sequential and contextual details of the agent’s interactions, creating a comprehensive narrative of its decision-making process (Line 15). The prompt design, inspired by Chain-of-Thought reasoning [47, 48], decomposes complex behavioral analysis into specific components. By incorporating detailed environmental parameters and quantitative evaluation guidelines, the prompt effectively guides the model’s analysis, enhancing its ability to generate insightful summaries of reinforcement learning behavior. This design was refined through iterative exploration, as detailed in I.
The prompt structure consists of four main components:
  • General Instructions: for example, “Generate a summary that captures the essence of the agent’s policy.”
  • Environment Description: such as, “Describe the agent’s task and key aspects of the environment.”
  • Interpretation Instructions: e.g., “Descriptions are from the agent’s perspective.”
  • Output Instructions: for instance, “Provide a concise summary.”
The LLM then produces the summary based on the prompt and the formatted experience dataset (Line 16).
Algorithm 1 SySLLM
Input: Environment $ENV$, trained policy $\pi$, captioners $C_{\text{observation}}$, $C_{\text{action}}$
Parameter: number of episodes $N$, prompt $P$
Output: summary $\in\mathcal{T}$
1: Initialize experience dataset EDED

### v3

@@
Abstract.
Explaining reinforcement learning agents is challenging because policies emerge from complex reward structures and neural representations that are difficult for humans to interpret. Existing approaches often rely on curated demonstrations that expose local behaviors but provide limited insight into an agent’s global strategy, leaving users to infer intent from raw observations. We propose SySLLM (Synthesized Summary using Large Language Models), a framework that reframes policy interpretation as a language-generation problem. Instead of visual demonstrations, SySLLM converts spatiotemporal trajectories into structured text and prompts an LLM to generate coherent summaries describing the agent’s goals, exploration style, and decision patterns. SySLLM scales to long-horizon, semantically rich environments without task-specific fine-tuning, leveraging LLM world knowledge and compositional reasoning to capture latent behavioral structure across policies. Expert evaluations show strong alignment with human analyses, and a large-scale user study found that 75.5% of participants preferred SySLLM summaries over state-of-the-art demonstration-based explanations. Together, these results position abstractive textual summarization as a paradigm for interpreting complex RL behavior. [^1] Code available at https://github.com/saharad1/SySLLM
Key words and phrases:
Reinforcement Learning, Policy Summarization, Explainable AI
{CCSXML}
¡ccs2012¿ ¡concept¿ ¡concept_id¿10010147.10010178.10010179.10010182¡/concept_id¿ ¡concept_desc¿Computing methodologies Natural language generation¡/concept_desc¿ ¡concept_significance¿500¡/concept_significance¿ ¡/concept¿ ¡concept¿ ¡concept_id¿10010147.10010257.10010258.10010261¡/concept_id¿ ¡concept_desc¿Computing methodologies Reinforcement learning¡/concept_desc¿ ¡concept_significance¿500¡/concept_significance¿ ¡/concept¿ ¡/ccs2012¿
1. Introduction
Reinforcement learning (RL) agents are increasingly deployed in sequential decision-making domains, yet their policies remain opaque to human stakeholders. This opacity limits trust, adoption, and effective debugging. While many explainable reinforcement learning (XRL) methods provide local insight into individual states or decisions, our focus is on global explanations that aim to capture an agent’s strategy across trajectories. Existing paradigms face trade-offs between expressiveness, scalability, and faithfulness. Extractive methods such as saliency maps greydanus2018visualizing; puri2019explain; atrey2019exploratory; samadi2024safe and demonstration-based policy summaries amir2018agent; frost2022explaining; liu2023learning; deshmukh2023explaining capture only fragments of behavior. Symbolic approaches such as rules or decision trees sequeira2020interestingness; peng2022inherently; mccarthy2022boolean; parham2023explaining often collapse in high-dimensional or partially observed environments. Across these approaches, a core flaw persists: they transform observations but do not synthesize them into holistic accounts, leaving users to reconstruct intent, adaptability, and failure modes from fragmented evidence anderson2020mental; amitaisurvey.
Large language models (LLMs) appear to offer a promising alternative. Their capacity for abstraction, compositional reasoning, and natural language generation zhao2023survey suggests that they could distill trajectories into human-readable accounts of agent behavior. However, applying them to RL policies is far from straightforward. LLMs are trained on static text-based corpora, while RL agents generate dynamic spatio-temporal trajectories grounded in states, actions, and rewards li2024advancing. This mismatch in modality makes faithful policy summarization non-trivial. Specifically, naive application risks producing fluent and plausibly sounding behavior descriptions that do not accurately reflect the underlying policy.
To address this challenge, we propose a textual-abstractive paradigm for policy summarization, in which explanations are expressed as natural language narratives integrating evidence from a database of policy execution traces, which we term the textual experience buffer (TEB). Unlike extractive or symbolic methods, this paradigm explicitly targets policy-level regularities and behavioral motifs. We formalize the task as a mapping from an agent’s experience buffer to a textual summary, guided by the desiderata of expressiveness, scalability, and faithfulness. To realize these principles in practice, we adopt a conceptual optimization view in which summaries balance coverage, parsimony, and fidelity. This formulation grounds both our methodology and evaluation, ensuring that system design choices are principled and directly address the unique challenges of applying LLMs to RL.
Building on this formulation, we introduce SySLLM, a framework that leverages LLMs to generate global policy summaries. SySLLM operates in two stages: agent–environment trajectories are first transformed into structured natural language descriptions of observations and actions, which are then synthesized into higher-level accounts through carefully designed prompting. To scale beyond context window limitations, SySLLM performs hierarchical summarization over large buffers. To mitigate variability in LLM outputs, it generates multiple candidate summaries and aggregates them into a consensus using embedding-based similarity. Together, these mechanisms ensure that SySLLM produces summaries that are general enough to capture policy-level regularities while specific enough to reflect distinctive behaviors.
We evaluated SySLLM across five MiniGrid environments and the Crafter domain, covering nine qualitatively distinct agent policies. Expert evaluation shows a strong alignment between SySLLM summaries and expert summaries, achieving high recall and precision scores that reflect faithful coverage of expert-identified behaviors. A user study with 192 participants further demonstrates that users strongly prefer textual summaries over demonstration-based summaries such as HIGHLIGHTS-DIV amir2018highlights, while performing equally well or better on policy identification tasks.
Our main contributions are threefold: (1) we introduce and formalize the task of abstractive-textual policy summarization in RL, framing it as a mapping from an agent’s experience buffer to natural language narratives defined by desiderata of expressiveness, scalability, and faithfulness; (2) we present SySLLM, a framework that leverages LLMs to synthesize structured trajectory descriptions into coherent global summaries, incorporating hierarchical summarization and consensus aggregation to address long horizons and variability in outputs, and (3) we provide extensive empirical validation through expert evaluations and a large-scale user study, demonstrating that SySLLM produces faithful summaries that are strongly preferred by participants over state-of-the-art demonstration-based baselines.
2. Related Work
Prior work in XRL spans both local explanations of individual decisions and global summaries of agent behavior, with the shared goal of improving policy interpretability for humans milani2022survey. Saliency and visualization methods highlight influential inputs greydanus2018visualizing; huber2022benchmarking but are often local and fragile; demonstration-based methods summarize behavior via selected trajectories amir2018agent; deshmukh2023explaining but place interpretive burden on users; surrogate models distill policies into rules or decision trees bastani2018verifiable; coppens2019distilling, though with fidelity–scalability trade-offs; and causal or reward-based explanations madumal2020explainable; juozapaitis2019explainable provide structured insights but typically require access to internals or domain expertise. Collectively, these techniques emphasize fragments or simplified proxies rather than synthesizing global accounts of an agent’s strategy.
Large language models (LLMs) bring complementary capabilities of abstraction, reasoning, and fluent text generation brown2020language. While prior work mainly uses LLMs during training—for example, guiding exploration or constructing world models du2023guiding; guan2023leveraging—their potential as explanation generators has received less attention. Some studies prompt LLMs to narrate behavior in real time wang2023voyager; yao2023react or to build symbolic simulators ahn2022can, but these efforts are often ungrounded in actual dynamics and focus on local rather than global explanations.
Natural language explanations for RL agents have also been studied. The early template-based methods hayes2017improving; peng2022inherently prioritized accessibility, but were brittle, while neural rationalization approaches translated trajectories into free form text ehsan2018rationalization; mccalmon2022caps. These provide interpretability but generally rely on handcrafted structures, focus on local justifications, or lack scalability across diverse scenarios. They rarely capture holistic behavioral patterns or leverage broader common sense knowledge. These limitations motivate our textual-abstractive paradigm, where LLMs synthesize coherent global summaries of agent behavior.
@@ 2. Related Work
Diagram of the data collection pipeline. The environment produces an observation grid showing objects like a yellow key and door. The observation is converted into a text description, then passed to an experience dataset. The agent’s policy network takes the observation and outputs an action, such as moving forward. The action and observation captions are stored in the dataset along with rewards.
Figure 1. Collecting the textual experience buffer (Section 4.2).
3. Problem Formulation
In this section, we formalize the abstractive-summarization problem. Specifically, we describe the setting, define the summarization task, and introduce conceptual principles that guide both our methodology and our evaluation.
3.1. Setting
We consider an RL environment modeled as a partially observable Markov decision process (POMDP):
$\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle,$
where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathcal{O}$ the observation space, $T(s^{\prime}\mid s,a)$ the transition kernel, $O(o\mid s,a)$ the observation function, $R(s,a)$ the reward function, and $\gamma\in(0,1]$ the discount factor.
An agent follows a stochastic policy
$\pi:\mathcal{O}\to\Delta(\mathcal{A}),$
which maps each observation $o\in\mathcal{O}$ to a distribution over actions. The interaction induces a distribution over trajectories
$\tau=(o_{0},a_{0},r_{0},\dots,o_{T}),\quad\tau\sim\pi,T,O.$
For summarization, we assume access not only to isolated trajectories but to an experience buffer
$\mathcal{B}_{\pi}=\{\tau_{1},\dots,\tau_{N}\},$
which aggregates multiple episodes sampled from $\pi$. This buffer serves as the raw material based on which textual explanations are generated.
3.2. Policy Summarization Task
We define a policy summarizer as a mapping
$f:\mathcal{B}_{\pi}\to\mathcal{T},$
where $\mathcal{T}$ denotes the space of abstractive textual explanations.
We require $f$ to approximate three key principles. Expressiveness: capture recurring behavioral patterns (e.g. “the agent prioritizes unlocking doors before exploring rooms”) rather than isolated actions. Scalability: operate over long horizons and large buffers while maintaining concise summaries. Faithfulness: reflect the actual distribution of behaviors under $\pi$, avoiding hallucinated or spurious strategies.
Conceptual Optimization View
We frame summarization as a conceptual optimization problem:
$T^{*}=\arg\max_{T\in\mathcal{T}}\;\mathcal{U}(T\mid\mathcal{B}_{\pi}),$
where $\mathcal{U}$ is a utility function balancing: coverage (operationalizing expressiveness by accounting for recurring behaviors), parsimony (operationalizing scalability through concise abstraction), and fidelity (operationalizing faithfulness via alignment with empirical evidence in $\mathcal{B}_{\pi}$).
This formulation is not solved directly. Instead, it provides a guiding lens for both methodology and evaluation: SySLLM instantiates coverage through multi-trajectory aggregation and hierarchical summarization, parsimony through representative summary selection, and fidelity through expert alignment and user validation. In the following sections, we detail how these principles are operationalized in practice and evaluated empirically.
@@ 3.2. Policy Summarization Task
SySLLM generating policy summaries
Figure 2. Generating global policy summaries (Section 4.3).
4. SySLLM Framework
Our SySLLM (Synthesized Summary using LLMs) framework formalizes policy summarization as a two-phase process: (i) experience collection and captioning, which transforms trajectories into natural language traces stored in a Textual Experience Buffer (TEB), and (ii) abstractive summarization, which synthesizes the TEB into a concise global description of the agent’s policy using a large language model (LLM). Algorithms 12 specify the pipeline, while Figures 12 illustrate its two phases.
4.1. Captioners
The first phase converts raw trajectories into textualized experiences. For a trajectory $\tau=(o_{1},a_{1},r_{1},\ldots,o_{T})$, we define two captioning functions:
$C_{\text{obs}}:\mathcal{O}\to\Sigma^{\star},\quad C_{\text{act}}:\mathcal{A}\to\Sigma^{\star},$
where $\Sigma^{\star}$ denotes the set of natural language strings. At each step $t$, the pair $(o_{t},a_{t})$ is mapped to $(C_{\text{obs}}(o_{t}),C_{\text{act}}(a_{t}))$, yielding a textual description of the observation and the action taken.
Observation Captioner. $C_{\text{obs}}$ produces structured descriptions of salient percepts (e.g. “The agent is facing a locked door with a key to the left.”).
Action Captioner. $C_{\text{act}}$ verbalizes the agent’s action (e.g. “move forward,” “pick up the key.”).
This follows prior work on language rounding jiang2019language; mirchandani2021ella. In practice, captioners may be rule-based, vision-language based, or hybrid; the concrete instantiation is left as a domain-specific choice and is discussed further in Section 9.
4.2. Constructing the Textual Experience Buffer
From the $N$ sampled episodes, we construct a Textual Experience Buffer (TEB):
$TEB_{\pi}=\langle e_{1},e_{2},\dots,e_{L}\rangle,\quad e_{t}=\langle\hat{o}_{t},\hat{a}_{t},r_{t},epID\rangle,$
where $\hat{o}_{t}=C_{\text{obs}}(o_{t})$ and $\hat{a}_{t}=C_{\text{act}}(a_{t})$ denote natural language captions of observations and actions. Each tuple records captioned observation, captioned action, reward, and episode identifier, preserving temporal coherence between steps.
Episodes are generated by sampling from the policy $\pi$:
$o_{1}\sim\texttt{RESET}(ENV),\quad a_{t}\sim\pi(o_{t}),\quad(o_{t+1},r_{t},done)\sim\texttt{STEP}(a_{t}),$
repeating until $done=\texttt{True}$. At each step, both observation and action are passed through captioning functions $C_{\text{obs}}$ and $C_{\text{act}}$.
The TEB is thus a textual analogue of a replay buffer: it aggregates multiple trajectories in a structured, language-based format that is directly consumable by LLMs, while maintaining the sequential structure necessary for policy-level reasoning. Further details on the fields stored in the TEB is provided in Appendix B.
4.3. Abstractive Summarization
Let $\mathcal{B}$ denote the space of textual experience buffers, where each $TEB_{\pi}\in\mathcal{B}$ is a finite ordered sequence of captioned experience tuples collected from executions of policy $\pi$.
The second phase maps the TEB to a global narrative $T\in\mathcal{T}$. We define the summarizer as
$f_{\theta}:\mathcal{B}\rightarrow\mathcal{T},\qquad T=f_{\theta}(TEB_{\pi}),$
where $\Sigma^{\star}$ denotes the space of finite token sequences and $f_{\theta}$ is instantiated by an LLM conditioned on a structured prompt.
Prompt Construction. The prompt follows a hierarchical structure, inspired by Chain-of-Thought reasoning wei2022chain; kojima2022large, which decomposes complex behavioral analysis into specific components. This design guides the LLM to progressively move from low-level traces in the TEB to higher-level abstractions of the policy. The prompt consists of:
  1. (1)
    General Instructions: define the summarization task.
  2. (2)
    Environment Context: describe task objectives and constraints.
  3. (3)
    Textual Experience Buffer: provide $TEB$ in structured form to preserve sequentiality.
  4. (4)
    Output Specification: constrain the output to a concise summary in natural language.
Prompt design details are provided in Appendix I.
Scalability via Hierarchical Summarization. A fundamental challenge in LLM-based summarization is the bounded context size: The TEB may exceed the maximum token budget $\kappa$ of the model. To address this, SySLLM employs a hierarchical procedure that recursively reduces the buffer until it fits within the context window.
Formally, define a summarization operator:
$\mathcal{S}_{\theta}:\mathcal{D}\to\Sigma^{\star},$
where $\mathcal{D}$ is any subset of the TEB and $\Sigma^{\star}$ the space of textual summaries. If $|TEB|\leq\kappa$, we directly apply:
$T=\mathcal{S}_{\theta}(TEB).$
If $|TEB|>\kappa$, we partition the buffer into $M$ disjoint subsets $TEB=\{TEB^{(1)},\dots,TEB^{(M)}\}$ such that each $|TEB^{(i)}|\leq\kappa$. For each subset we compute intermediate summaries:
$S_{i}=\mathcal{S}_{\theta}(TEB^{(i)}),\quad i=1,\dots,M,$
and aggregate them by applying $\mathcal{S}_{\theta}$ again:
$T=\mathcal{S}_{\theta}(\{S_{1},\dots,S_{M}\}).$
This recursive divide-and-conquer scheme:
$T=\textsc{HierarchicalSummarize}(TEB,\kappa)$
ensures that SySLLM remains applicable to arbitrarily large buffers while preserving coverage across all episodes and compressing details into intermediate summaries.
Candidate Generation and Selection. Once the input (original or hierarchical) fits within context, we query the LLM to generate $K$ candidate summaries $\{T_{1},T_{2},\dots,T_{K}\}$ via stochastic decoding. This captures variability in abstraction and phrasing. To select a robust final summary, each candidate is embedded into a semantic vector space using a pretrained embedding model, yielding $\phi(T_{i})\in\mathbb{R}^{d}$. We compute the centroid:
$c=\tfrac{1}{K}\sum_{i=1}^{K}\phi(T_{i}),$
and measure distances:
$d_{i}=\lVert\phi(T_{i})-c\rVert_{2}.$
The final summary $T^{*}$ is chosen as the median representative, i.e. the candidate closest to the median-ranked distance from the centroid:
$T^{*}=\arg\min_{T_{i}}\;\big|\text{rank}(d_{i})-\tfrac{K}{2}\big|.$
This selection scheme balances generality (summaries near the centroid capture broad regularities) with specificity (summaries farther away capture contextual details), yielding a consensus-style narrative.
Algorithm 1 SySLLM Framework
Input: Environment $ENV$, trained policy $\pi$, captioners $C_{\text{obs}}$, $C_{\text{act}}$, base prompt $P$
Parameters: number of episodes $N$, token budget $\kappa$, #candidates $K$, embedding model $\phi(\cdot)$
Output: Policy summary $T^{*}\in\mathcal{T}$
1:Initialize Textual Experience Buffer $TEB\leftarrow\varnothing$